# Rethinking Improved Privacy-Utility Trade-off with Pre-existing Knowledge for DP Training

### v1

@@ Abstract
Differential privacy (DP) provides a provable framework for protecting individuals by customizing a random mechanism over a privacy-sensitive dataset. Deep learning models have demonstrated privacy risks in model exposure as an established learning model unintentionally records membership-level privacy leakage. Differentially private stochastic gradient descent (DP-SGD) has been proposed to safeguard training individuals by adding random Gaussian noise to gradient updates in the backpropagation. Researchers identify that DP-SGD typically causes utility loss since the injected homogeneous noise alters the gradient updates calculated at each iteration. Namely, all elements in the gradient are contaminated regardless of their importance in updating model parameters. In this work, we argue that the utility loss mainly results from the homogeneity of injected noise. Consequently, we propose a generic differential privacy framework with heterogeneous noise ($\mathsf{DP\text{-}Hero}$) by defining a heterogeneous random mechanism to abstract its property. The insight of $\mathsf{DP\text{-}Hero}$ is to leverage the knowledge encoded in the previously trained model to guide the subsequent allocation of noise heterogeneity, thereby leveraging the statistical perturbation and achieving enhanced utility. Atop $\mathsf{DP\text{-}Hero}$, we instantiate a heterogeneous version of DP-SGD, where the noise injected into gradients is heterogeneous and guided by prior-established model parameters. We conduct comprehensive experiments to verify and explain the effectiveness of the proposed $\mathsf{DP\text{-}Hero}$, showing improved training accuracy compared with state-of-the-art works. Broadly, we shed light on improving the privacy-utility space by learning the noise guidance from the pre-existing leaked knowledge encoded in the previously trained model, showing a different perspective of understanding the utility-improved DP training.
Index Terms:
Differential Privacy, Heterogeneous Noise, DP-SGD, Privacy-Utility Trade-off.
I Introduction
@@
Deep learning has achieved remarkable success across a wide spectrum of domains [1, 2, 3, 4, 5, 6], primarily due to the massive data used for model training. As training data has been thoroughly analyzed to optimize model performance, a significant privacy concern arises regarding the model’s potential to memorize individual data points [7, 8, 9]. Indeed, numerous existing studies [10, 11, 12] have demonstrated the feasibility of identifying the presence of particular records or verbatim texts in the training dataset, thereby raising severe privacy concerns.
@@
Differential privacy (DP) [13, 14, 15, 16], emerged as de facto protection, can provide provable security for individuals’ privacy by adding the i.i.d noise to the sensitive data or computations. In detail, DP guarantees statistical indistinguishability between the outputs of two random mechanisms, which originate from the private inputs with or without a substituted individual data point. To protect sensitive data used in the training process, differentially private stochastic gradient descent (DP-SGD) [14] has been proposed and regarded as a main-steam method. The idea of DP-SGD is to add the homogeneous noise sampled from a Gaussian distribution to the aggregated gradient derived from a batch of examples in every training iteration. Accordingly, DP-SGD can thwart an adversary from attaining membership of a particular data memorized by model parameters when the adversary dissects an established model. One could adopt DP-SGD as a baseline [17, 18] for supporting secure non-convex training for neural networks.
@@
Subsequently, researchers identify the inherent trade-off between privacy and utility introduced by DP-SGD. It is a well-known challenge to achieve high model utility/performance given meaningful DP guarantees [19, 20, 21, 16, 22, 23] since acquiring strong protection realized by a large noise scale generally leads to unavoidable utility loss and performance degrading. For example, the number of DP-SGD training iterations may increase by 10×10$\times$10 × towards a similar utility metric compared with the pure SGD. Accordingly, a research line of works [20, 21, 16, 22] explored to acquire a better utility by flexibly and empirically calibrate privacy budget allocation. Regarding composition theorem, they try to either reallocate/optimize the privacy budget [20, 16, 22, 23, 24] or modify the clip-norms [25, 26] of a (set of) fixed noise distribution(s) in each iteration. These dynamic noise allocation solutions optimize the noise allocation in the range of the whole training process with a constant budget. Sharing the same spirit as DP-SGD, these methods employ homogeneous noise to perturb gradient updates.
@@ I-A Technical Overview
1) Allocating noise heterogeneity via pre-existing knowledge. To generate the model-guided heterogeneity, we propose a novel dynamic noise allocation scheme, where an iteration-wise (for short, stateful) matrix $\mathsf{SVec}^{(t-1)}$ is computed using the pre-existing model at $(t-1)$-th iteration. With the notion of stateful $\mathsf{SVec}^{(t-1)}$, we can guide the noise heterogeneity at the $t$-th training iteration. Namely, the stateful $\mathsf{SVec}$ adjusts the noise $n$ used to perturb gradient updates at every iteration according to the heterogeneity derived by the $\mathsf{SVec}^{(t-1)}$. Consequently, the posterior random mechanism is guided by pre-existing knowledge encoded in prior model parameters at every training iteration. Specifically, we formally define our novel scheme $\mathsf{DP\text{-}Hero}$ as a random mechanism that $\mathcal{M}^{(t)}=\mathbf{g}^{(t)}+\mathsf{SVec}^{(t-1)}\cdot\mathbf{n}$, where the abstraction of $\mathsf{SVec}^{(t-1)}$ is independent to knowledge extraction function $\mathcal{F}$ of learned model $\mathsf{NNet}$ and indexed by states $t-1,t$.
For theoretical analysis, we abstract the notion of heterogeneous DP learning with stateful guidance for allocating noise heterogeneity. By adopting composition [27, 28] and Rényi Divergence, we provide theoretical analysis on $\mathsf{DP\text{-}Hero}$ SGD training following conventional proof style. Accordingly, the instantiation of $\mathsf{DP\text{-}Hero}$ SGD, regarded as a modified variant of standard DP-SGD, attains the standard DP guarantee.
2) Constructing heterogeneous DP-SGD. We instantiate $\mathsf{DP\text{-}Hero}$ as a heterogeneous version of DP-SGD, where the noise injected into gradient updates is heterogeneous. Specifically, the stateful $\mathsf{SVec}^{(t-1)}$ at the $(t-1)$-th training iteration is derived from decomposition on model parameters $\mathbf{W}^{(t-1)}$ at the prior training iteration, capturing the pre-existing knowledge. Knowledge involved in $\mathbf{W}^{(t-1)}$, serving as allocation guidance, propagates to the DP noise injected to gradients at the $t$-th training iteration, following the style of DP-SGD. Accordingly, it captures the pre-existing statistical knowledge of private training data, extracting heterogeneity applied to features. Later, the stateful guidance matrix $\mathsf{SVec}^{(t-1)}$ adjusts the parameters of Gaussian distribution, equivalently affecting the heterogeneity of noise added to diverse features in the back-propagation of SGD. Prior knowledge from extracted features has been reasonably DP-protected, thus not incurring extra risks in exposing private data. The plug-in $\mathsf{DP\text{-}Hero}$ SGD is generic and independent of learning models, best for both worlds in performance and privacy.
@@ I-B Contribution Summary
II Related Works
II-A Differential Privacy for Deep Learning
Differential privacy has emerged as a solid solution to safeguard privacy in the field of deep learning. Differential privacy (DP) for deep learning can be classified into four directions: input perturbation [29, 30], output perturbation [15], objective perturbation [31, 32], and utility optimization [33, 14, 34, 16], showing valuable insights in the aspects of theory and practice. DP could quantify to what extent individual privacy (i.e., whether a data point contributes to the training model) in a statistical dataset is preserved while releasing the established model trained over the specific dataset. Typically, DP learning has been accomplished by applying the unbiased Gaussian noise to the gradient descent updates, a notable example of DP-SGD [14]. To be specific, DP-SGD adds the i.i.d. noise sampled from Gaussian distribution to model gradients to protect example-level training data involved in the training process in every iteration.
The noise-adding mechanism has been widely adopted in various learning algorithms, e.g., image classification and natural language processing. PATE [15] is an approach to providing differentially private aggregation of a teacher-student model. Due to the property of post-processing [35], the student’s model is differentially private since it trains over the noisy inputs. Bayesian differential privacy [36] takes into account the data distribution for practicality [37]. By instantiating hypothetical adversaries [38], various threat models are employed to show corresponding levels of privacy leakage from both the views of practitioners and theoreticians.
Privacy auditions and robustness, or cryptographic protection [39, 40] belong to orthogonal research directions, focusing on the evaluative estimation of the privacy guarantee or cipher-text transmission. Membership inference attack [41] enables detecting the presence or absence of an individual example, implying a lower bound on the privacy parameter $\epsilon$ via statistics [42]. Notably, Steinke et al. [43] theoretically proves the feasibility of auditing privacy through membership inference on multiple examples simultaneously, elaborating an efficient one-round solution. Combining different techniques with this work can be promising, while it is out of scope for this work.
II-B Privacy-Utility Trade-off
For acquiring higher utility [44], recent works explore the adaptive mechanism of DP training from different perspectives. They try to either reallocate/optimize the privacy budget [20, 21, 16, 22, 23, 45] or modify the clip-norms [25, 26] of a (set of) fixed noise distribution(s) in each iteration. Such a branch of work points out a promising direction of adaptivity via redesigning the randomness. Privacy budget scheduling [23] improves the utility of differentially private algorithms in various scenarios. Unlike the aforementioned advances of dynamic noise allocation, our exploration of adjusting noise heterogeneity by model parameters aims to improve the utility of the established model at every iteration rather than optimizing the noise allocation in the range of the whole training process with a constant budget. Handcrafted features, learned from public data, can improve model utility given a fixed privacy budget [46]. Rather than introducing auxiliary data, we aim to extract knowledge from protected model parameters without extra data assistance.
Previous analyses have enabled an understanding of utility bounds for DP-SGD mainly in an empirical manner. Altschuler and Talwar [47] explored the theory foundation of privacy loss – how sensitive the output of DP-SGD is. They solve a tighter utility bound given the privacy loss as a function of the number of iterations, concluding that after a small burn-in period, running DP-SGD longer leaks no further privacy. In this work, we exploit visual explanation [48] and theoretical understanding to explore the essence of privacy-utility space.
III Preliminary
III-A General Notion of Differential Privacy
Differential privacy (DP) [13, 35] theoretically guarantees individual privacy that the algorithm’s output changes insignificantly (see Definition 2) if the inputting data has small perturbations. Pure $\epsilon$-differential privacy is difficult to achieve in realistic learning settings, whereas the seminal work [14] training with SGD adopts approximate ($\epsilon,\delta$)-differential privacy, formally defined below.
@@ III-A General Notion of Differential Privacy
[…]
Next, we introduce the definition of privacy loss [35] on an outcome $o$ as a random variable when DP operates on two adjacent databases $D$ and $D^{\prime}$. Privacy loss is a random variable that accumulates the random noise added to the algorithm/model.
Definition 3 (Privacy Loss [35]).
@@
III-B DP Stochastic Gradient Descent
DP-SGD [14], regarded as a landmark work, has been proposed to safeguard example-level model knowledge encoded from the training data, constrained by the privacy budget allocated at each training iteration. As reported by DP-SGD, adding i.i.d. noise inevitably brings parameter perturbation over the learned model in practice. Research efforts such as [19, 20, 21, 16, 22, 23] are focused on developing techniques that can provide stronger privacy guarantees while minimizing the loss of utility from various perspectives, for example, clipping value optimization and privacy budget crafting.
@@ III-B DP Stochastic Gradient Descent
Through revisiting DP-SGD, we explore explaining the root of utility loss and bridge the concept of model-knowledge guidance and DP, making a DP training process fitting to enhance privacy-utility trade-offs better. We showcase new thinking – not employing auxiliary (e.g., public data) assistance for the higher model, and thus rethinking tolerant leakage (statistical knowledge, not membership, aligning standard DP definition) encoded in the prior DP-trained model.
III-C Rényi Differential Privacy
Rényi differential privacy [28] has been proposed as a natural relaxation of differential privacy, particularly suitable for composing privacy guarantee of heterogeneous random mechanisms derived from algorithms. zCDP [27] and Rényi DP [28] (RDP) are defined through Rényi Divergence by Bun et al. [27] for a tight analysis, thereby providing accumulating cumulative loss accurately and strong privacy guarantees. Definition 4 presents the Rényi Divergence [28] for defining the Rényi differential privacy [28] as Definition 5.
Definition 4 (Rényi Divergence [28]).
@@ III-C Rényi Differential Privacy
Definition 5 (Rényi Differential Privacy [28]).
@@
Theorem 1 (From RDP to $(\epsilon,\delta)$-DP [28]).
If $\mathcal{M}$ is an $(\alpha,\epsilon)$-RDP mechanism, it also satisfies $(\epsilon+\frac{\log 1/\delta}{\alpha-1},\delta)$-DP for any $10<\delta<$ 1.
III-D Security Model for Centralized DP Learning
@@ IV Noise Heterogeneity in DP
IV-A Define Heterogeneous DP Learning
We rethink reasonable leakages in DP models and make use of the pre-existing knowledge learned in the current model parameters to improve subsequent DP training performance. Model training starts with a random $\mathsf{NNet}^{(0)}$ towards a convergent model $\mathsf{NNet}^{(T)}$, which captures knowledge learned from data iteration by iteration. Naturally, our idea is to introduce a scalar vector $\mathsf{SVec}$ that is derived from the learned knowledge in $\mathsf{NNet}$ in the prior training process to serve as the guidance for subsequent DP training.
Consider a function $\mathcal{F}$ to denote functionality achieved by neural networks $\mathsf{NNet}$. The $\mathsf{NNet}^{(t)}$, trained with the DP mechanism, denotes the deep learning model at iteration $t$. We formulate DP trained model at $t$-th iteration to be $\mathcal{F}(\mathsf{NNet}^{(t)},\mathsf{Data})$ given private $\mathsf{Data}$. We utilize the $\mathsf{SVec}^{(t-1)}$ at $t-1$-th iteration to adjust the next-step noise allocation at $t$-th iteration, where $\mathsf{SVec}^{(t-1)}$ is computed by the prior $\mathsf{NNet}^{(t-1)}$ at $(t-1)$-th iteration involving features learned in the first $t-1$ iterations. Concretely, Definition 6 introduces a general notion of heterogeneous DP learning ($\mathsf{DP\text{-}Hero}$) that varies with the time $t$, essentially adjusting the noise vector (sampled from Gaussian distribution) operated over the learning model.
@@ IV-A Define Heterogeneous DP Learning
$\mathcal{N}(\mu,\sigma^{2})$ represents noise distribution with parameters $\mu,\sigma$. To generate pre-existing knowledge stored in the model parameters, we can employ a knowledge-extraction method (e.g., principal component analysis [49]) to extract pre-existing knowledge stored in the learned model, saying $\mathsf{SVec}^{(t-1)}\propto\mathcal{F}_{\mathsf{know}}(\mathsf{NNet}^{(t-1)}),t\in[0,T]$. Accordingly, the noise sampled from the Gaussian distribution is scaled by $\mathsf{SVec}$ (i.e., values and noise direction). The $\mathsf{SVec}$ keeps varied for tracking DP model training, calibrating noise vector via pre-existing knowledge stored in the model. In summary, the $\mathsf{DP\text{-}Hero}$ expects to: 1) be tailored with heterogeneous DP noise that is added to the learning process; 2) be generic and irrelevant to the convergence route for distinctive models for iteratively reaching a model optimum; 3) have good model accuracy and convergence performance given a preset privacy budget.
@@
    ((((Privacy)))). If for any two neighboring datasets ${D}$ and ${D}^{\prime}$, $\mathrm{Pr}(\mathcal{M}^{(t)}({D})\in S)\leq e^{\epsilon_{t}}\cdot\mathrm{Pr}(\mathcal{M}^{(t)}({D}^{\prime})\in S)+\delta_{t}$ for any iteration $t\in[0,T]$.
  • ii
    ((((Utility)))). Supposing an optimal $Z^{\ast}$, the objective function satisfies $\arg\ \min_{\epsilon_{t}=\epsilon/T}\mathcal{F}_{\mathsf{Diff}}[\mathcal{M}^{(t)}\|Z^{\ast}]$.
@@
    $(t$-Sequential Composition)))). If ~$\tilde{\mathcal{M}}=(\mathcal{M}^{(0)},\ldots,\mathcal{M}^{(t)},\allowbreak\ldots,\mathcal{M}^{(T)})$, $\tilde{\mathcal{M}}$ satisfies $(\tilde{\epsilon},\delta)$-DP such that $\tilde{\epsilon}\leq\epsilon$.
Property (i) essentially guarantees differential privacy [13, 35] at each training iteration. Property (ii) extracts the iteration-wise optimization, which expects that the difference measurement $\mathcal{F}_{\mathsf{Diff}}$ between the noisy model and pure model are small as possible. Given a fixed privacy budget $\epsilon/T$, improving utility expects to reduce the difference between $\mathcal{M}^{(t)}$ and non-noisy $Z^{\ast}$. Property (ii) asks for no extra privacy leakages in the $\mathsf{DP\text{-}Hero}$ under privacy composition, which is the same as the standard DP guarantee.
IV-B Overview of DP Heterogeneous SGD
Before constructing DP heterogeneous SGD ($\mathsf{DP\text{-}Hero}$ SGD), we adopt the notations of DP-SGD by revisiting standard DP-SGD [14]. DP-SGD trains a model with parameters $\mathbf{W}$ by minimizing the empirical loss function $\mathcal{L}(\mathbf{W})$. For a random example $x_{i}$, DP-SGD computes gradients $\mathbf{g}(x_{i})\leftarrow\nabla(\mathbf{W},x_{i})$ with clipping value $C$, and adds noise to $\sum_{i}\mathbf{g}(x_{i})$ sampled from Gaussian distribution $\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})$. An adversary cannot view the training process except for the DP-trained model.

### v3

# Revisiting Privacy-Utility Trade-off for DP Training with Pre-existing Knowledge

@@ Abstract
Differential privacy (DP) provides a provable framework for protecting individuals by customizing a random mechanism over a privacy-sensitive dataset. Deep learning models have demonstrated privacy risks in model exposure as an established learning model unintentionally records membership-level privacy leakage. Differentially private stochastic gradient descent (DP-SGD) has been proposed to protect training individuals by adding random Gaussian noise to gradient updates in backpropagation. Researchers identified that DP-SGD causes utility loss, as the homogeneous noise injected can alter the gradient updates calculated at each iteration. Namely, all elements in the gradient are contaminated regardless of their importance in updating model parameters. In this work, we argue that the utility can be optimized by involving the heterogeneity of the injected noise. Consequently, we propose a generic differential privacy framework with heterogeneous noise ($\mathsf{DP\text{-}Hero}$) by defining a heterogeneous random mechanism to abstract its property. The insight of $\mathsf{DP\text{-}Hero}$ is to leverage the knowledge encoded in the previously trained model to guide the subsequent allocation of noise heterogeneity, thereby leveraging the statistical perturbation and achieving enhanced utility. Atop $\mathsf{DP\text{-}Hero}$, we instantiate a heterogeneous version of DP-SGD and further extend it to federated training. We conduct comprehensive experiments to verify and explain the effectiveness of the proposed $\mathsf{DP\text{-}Hero}$, showing improved training accuracy compared with state-of-the-art works. Broadly, we shed light on improving the privacy-utility space by learning the noise guidance from the pre-existing leaked knowledge encoded in the previously trained model, showing a different perspective of understanding the utility-improved DP training.
Introduction
@@
Deep learning has achieved remarkable success across a wide spectrum of domains (Lu et al. 2022; Zhang et al. 2022; Viswanathan et al. 2023; Zhao et al. 2024), primarily relying on the massive data used for model training. As training data has been thoroughly analyzed to optimize model performance, a significant privacy concern arises regarding the model’s potential to memorize individual data points (Biderman et al. 2023; Carlini et al. 2023; Lukas et al. 2023; Zheng et al. 2023). Indeed, a growing body of studies (Shokri et al. 2017; Hitaj, Ateniese, and Pérez-Cruz 2017; Carlini et al. 2021) have demonstrated that it is feasible to identify the presence of a particular record or verbatim texts in the training dataset, thereby raising severe privacy concerns.
@@
Differential privacy (DP) (Dwork et al. 2006; Abadi et al. 2016; Papernot et al. 2018; Yu et al. 2019), emerged as de facto protection, can provide provable security for individuals’ privacy by adding the i.i.d noise to the sensitive data or computations. In detail, DP guarantees statistical indistinguishability between the outputs of two random mechanisms, which originate from the private inputs with or without a substituted individual data point. To protect sensitive data used in the training process, differentially private stochastic gradient descent (DP-SGD) (Abadi et al. 2016) has been proposed and regarded as a main-steam method. The idea of DP-SGD is to add the homogeneous noise sampled from a Gaussian distribution to the aggregated gradient derived from a batch of examples in every training iteration. Accordingly, DP-SGD, serving as the most popular baseline, can thwart an adversary from attaining membership of a particular data record when the adversary dissects an established model.
@@
Subsequently, researchers identified the inherent trade-off between privacy and utility introduced by DP-SGD. It is a well-known challenge to achieve high model utility/performance given meaningful DP guarantees (Tang et al. 2023; Lee and Kifer 2018; Mohammady et al. 2020; Yu et al. 2019; Geng and Viswanath 2016; Luo et al. 2021) since acquiring strong protection realized by a large noise scale generally leads to unavoidable utility loss and performance degrading. For example, the number of DP-SGD training iterations may increase by $10\times$ towards a similar utility metric compared with the pure SGD. Accordingly, a research line of works (Lee and Kifer 2018; Mohammady et al. 2020; Yu et al. 2019; Geng and Viswanath 2016) explored to acquire a better utility by flexibly and empirically calibrate privacy budget allocation. Regarding composition theorem, they try to either reallocate/optimize the privacy budget (Lee and Kifer 2018; Yu et al. 2019; Geng and Viswanath 2016; Luo et al. 2021; Yang et al. 2023) or modify the clip-norms (Pichapati et al. 2019; van der Veen et al. 2018) of a (set of) fixed noise distribution(s) in each iteration. These dynamic noise allocation solutions optimize the noise allocation in the range of the whole training process with a preset budget, but employs homogeneous noise at each iteration to perturb gradient updates.
@@
1) Allocating noise heterogeneity via pre-existing knowledge. To generate the model-guided heterogeneity, we propose a novel dynamic noise allocation scheme, where an iteration-wise (for short, stateful) matrix $\mathsf{SVec}^{(t-1)}$ is computed using the pre-existing model at (t1)(t-1)-th iteration. With the notion of stateful $\mathsf{SVec}^{(t-1)}$, we can guide the noise heterogeneity at the $t$-th training iteration. Namely, the stateful $\mathsf{SVec}$ adjusts the noise $n$ used to perturb gradient updates at every iteration according to the heterogeneity derived by the $\mathsf{SVec}^{(t-1)}$. Consequently, the posterior random mechanism is guided by pre-existing knowledge encoded in prior model parameters at every training iteration. Specifically, we consider a random mechanism named as $\mathsf{DP\text{-}Hero}$ that adds heterogeneous noise $\mathsf{SVec}^{(t-1)}\cdot\mathbf{n}$ to gradients $\mathbf{g}^{(t)}$: $\mathcal{M}^{(t)}=\mathbf{g}^{(t)}+\mathsf{SVec}^{(t-1)}\cdot\mathbf{n}$, where the abstraction of $\mathsf{SVec}^{(t-1)}$ is independent to knowledge extraction function $\mathcal{F}$ of learned model $\mathsf{NNet}$ and indexed by states t1,tt-1,t.
For theoretical analysis, we abstract the notion of heterogeneous DP learning with stateful guidance for allocating noise heterogeneity. By adopting composition (Bun and Steinke 2016; Mironov 2017) and Rényi Divergence, we provide theoretical analysis on $\mathsf{DP\text{-}Hero}$ SGD training following conventional proof style. Accordingly, the instantiation of $\mathsf{DP\text{-}Hero}$ SGD, regarded as a modified variant of standard DP-SGD, attains the standard DP guarantee.
2) Constructing heterogeneous DP-SGD. We instantiate $\mathsf{DP\text{-}Hero}$ as a heterogeneous version of DP-SGD, where the noise injected into gradient updates is heterogeneous. Specifically, the stateful $\mathsf{SVec}^{(t-1)}$ at the (t1)(t-1)-th training iteration is derived from decomposition on model parameters $\mathbf{W}^{(t-1)}$ at the prior training iteration, capturing the pre-existing knowledge. Knowledge involved in $\mathbf{W}^{(t-1)}$, serving as allocation guidance, propagates to the DP noise injected to gradients at the $t$-th training iteration, following the style of DP-SGD. Accordingly, it captures the pre-existing statistical knowledge of private training data, extracting heterogeneity applied to features. Later, the stateful guidance matrix $\mathsf{SVec}^{(t-1)}$ adjusts the parameters of Gaussian distribution, equivalently affecting the heterogeneity of noise added to diverse features in the back-propagation of SGD. Prior knowledge from extracted features has been reasonably DP-protected, thus not incurring extra risks in exposing private data. The plug-in $\mathsf{DP\text{-}Hero}$ SGD is generic and independent of learning models, best for both worlds in performance and privacy. To demonstrate the generalization of $\mathsf{DP\text{-}Hero}$, we present an extension of FedFed (Yang et al. 2023) by replacing its SGD optimizer with $\mathsf{DP\text{-}Hero}$ SGD. This modification enables FedFed to achieve rigorous DP guarantees for client data while retaining its scalability and communication efficiency.
@@
    To apply $\mathsf{DP\text{-}Hero}$ SGD, we present an extension of FedFed (Yang et al. 2023) with formal DP guarantee in federated learning.
  4. 4.
@@
Preliminary
General Notion of Differential Privacy
Differential privacy (DP) (Dwork et al. 2006; Dwork and Roth 2014) theoretically guarantees individual privacy that the algorithm’s output changes insignificantly (see Definition 2) if the inputting data has small perturbations. Pure $\epsilon$-differential privacy is difficult to achieve in realistic learning settings, whereas the seminal work (Abadi et al. 2016) training with SGD adopts approximate ($\epsilon,\delta$)-differential privacy, formally defined below.
@@
The sensitivity of a query function $\mathcal{F}:\mathbb{D}\rightarrow\mathbb{R}$ for any two neighboring datasets $D,D^{\prime}$ is,
$\Delta=\max_{D,D^{\prime}}\|\mathcal{F}(D)-\mathcal{F}(D^{\prime})\|,$ (2)
where $\|\cdot\|$ denotes $L_{1}$ or $L_{2}$ norm.
Next, we introduce the definition of privacy loss (Dwork and Roth 2014) on an outcome $o$ as a random variable when DP operates on two adjacent databases $D$ and $D^{\prime}$. Privacy loss is a random variable that accumulates the random noise added to the algorithm/model.
Definition 3 (Privacy Loss (Dwork and Roth 2014)).
@@
DP Stochastic Gradient Descent
DP-SGD (Abadi et al. 2016), regarded as a landmark work, has been proposed to safeguard example-level model knowledge encoded from the training data, constrained by the privacy budget allocated at each training iteration. As reported by DP-SGD, adding i.i.d. noise inevitably brings parameter perturbation over the learned model in practice. Research efforts such as (Tang et al. 2023; Lee and Kifer 2018; Mohammady et al. 2020; Yu et al. 2019; Geng and Viswanath 2016; Luo et al. 2021) are focused on developing techniques that can provide stronger privacy guarantees while minimizing the loss of utility from various perspectives, e.g., clipping value optimization and privacy budget crafting. Zhou et al. (Zhou, Wu, and Banerjee 2021) improves utility by projecting the noisy gradients to a low-dimensional subspace, while $\mathsf{DP\text{-}Hero}$ explores reusing the DP protected statistical knowledge learned from the private data.
@@
Through revisiting DP-SGD, we explore explaining the root of utility loss and bridge the concept of model-knowledge guidance and DP, making a DP training process fitting to enhance privacy-utility trade-offs better. We showcase new thinking – not employing auxiliary (e.g., public data) assistance for the higher model utility, and thus rethinking tolerant leakage (statistical knowledge, not membership, aligning standard DP definition) encoded in the prior DP-trained model.
Rényi Differential Privacy
Rényi differential privacy (Mironov 2017) has been proposed as a natural relaxation of differential privacy, particularly suitable for composing privacy guarantee of heterogeneous random mechanisms derived from algorithms. zCDP (Bun and Steinke 2016) and Rényi DP (Mironov 2017) (RDP) are defined through Rényi Divergence by Bun et al. (Bun and Steinke 2016) for a tight analysis, thereby providing accumulating cumulative loss accurately and strong privacy guarantees. Definition 4 presents the Rényi Divergence (Mironov 2017) for defining the Rényi differential privacy (Mironov 2017) as Definition 5.
Definition 4 (Rényi Divergence (Mironov 2017)).
@@
Definition 5 (Rényi Differential Privacy (Mironov 2017)).
@@
Theorem 1 (From RDP to $(\epsilon,\delta)$-DP (Mironov 2017)).
If $\mathcal{M}$ is an $(\alpha,\epsilon)$-RDP mechanism, it also satisfies $(\epsilon+\frac{\log 1/\delta}{\alpha-1},\delta)$-DP for any $0<\delta<1$.
Security Model for Centralized DP Learning
@@
Heterogeneity Guidance
Adding homogeneous DP noise uniformly across all model parameters often leads to suboptimal utility due to a mismatch between the injected noise magnitude and the underlying parameter sensitivity. Model parameters typically exhibit substantial heterogeneity in their magnitudes, gradient sensitivities, and contributions to overall task loss; applying the identical noise scale irrespective of these factors disproportionately corrupts low-norm or highly sensitive parameters, severely reducing their effective signal-to-noise ratio. Conversely, larger or less-sensitive parameters can tolerate greater perturbations without substantial impact on convergence. Heterogeneous noise schemes address this issue by varying the injected noise based on structural properties such as parameter norm scaling, per-layer sensitivity, or gradient clipping statistics. By preserving a more uniform effective SNR across parameters, heterogeneous noise enables models to maintain higher expressive power and faster convergence under the same global DP guarantee, yielding significantly improved privacy-utility tradeoffs.
We revisit reasonable leakages in DP models and make use of the pre-existing knowledge learned in the current model parameters to improve subsequent DP training performance. Model training starts with an initial random model $\mathsf{NNet}^{(0)}$ towards a convergent model $\mathsf{NNet}^{(T)}$, which captures knowledge learned from data iteration by iteration. Naturally, our idea is to introduce a scalar vector $\mathsf{SVec}$ that is derived from the learned knowledge in $\mathsf{NNet}$ in the prior training process to serve as the guidance for subsequent DP training.
Define Heterogeneous DP Learning
Consider a function $\mathcal{F}$ to denote functionality achieved by neural networks $\mathsf{NNet}$. The $\mathsf{NNet}^{(t)}$, trained with the DP mechanism, denotes the deep learning model at iteration $t$. We formulate DP trained model at $t$-th iteration to be $\mathcal{F}(\mathsf{NNet}^{(t)},\mathsf{Data})$ given private $\mathsf{Data}$. We utilize the $\mathsf{SVec}^{(t-1)}$ at t1t-1-th iteration to adjust the next-step noise allocation at $t$-th iteration, where $\mathsf{SVec}^{(t-1)}$ is computed by the prior $\mathsf{NNet}^{(t-1)}$ at (t1)(t-1)-th iteration involving features learned in the first t1t-1 iterations. Concretely, Definition 6 introduces a general notion of heterogeneous DP learning ($\mathsf{DP\text{-}Hero}$) that varies with the time $t$, essentially adjusting the noise vector (sampled from Gaussian distribution) operated over the learning model.
@@
$\mathcal{N}(\mu,\sigma^{2})$ represents noise distribution with parameters $\mu,\sigma$. To generate pre-existing knowledge stored in the model parameters, we can employ a knowledge-extraction method (e.g., principal component analysis (Jolliffe and Cadima 2016)) to extract pre-existing knowledge stored in the learned model, saying $\mathsf{SVec}^{(t-1)}\propto\mathcal{F}_{\mathsf{know}}(\mathsf{NNet}^{(t-1)}),t\in[0,T]$. Accordingly, the noise sampled from the Gaussian distribution is scaled by $\mathsf{SVec}$ (i.e., values and noise direction). The $\mathsf{SVec}$ keeps varied for tracking DP model training, calibrating noise vector via pre-existing knowledge stored in the model. In summary, the $\mathsf{DP\text{-}Hero}$ expects to: 1) be tailored with heterogeneous DP noise that is added to the learning process; 2) be generic and irrelevant to the convergence route for distinctive models for iteratively reaching a model optimum; 3) have good model accuracy and convergence performance given a preset privacy budget.
@@
    ((Privacy)). If for any two neighboring datasets D{D} and ${D}^{\prime}$, $\mathrm{Pr}(\mathcal{M}^{(t)}({D})\in S)\leq e^{\epsilon_{t}}\cdot\mathrm{Pr}(\mathcal{M}^{(t)}({D}^{\prime})\in S)+\delta_{t}$ for any iteration $t\in[0,T]$.
  • $i$
    ((Utility)). Supposing an optimal $Z^{\ast}$, the objective function satisfies $\arg\ \min_{\epsilon_{t}=\epsilon/T}\mathcal{F}_{\mathsf{Diff}}[\mathcal{M}^{(t)}\|Z^{\ast}]$.
@@
    $(t$-Sequential Composition)). If $\tilde{\mathcal{M}}=(\mathcal{M}^{(0)},\ldots,\mathcal{M}^{(t)},\allowbreak\ldots,\mathcal{M}^{(T)})$, $\tilde{\mathcal{M}}$ satisfies $(\tilde{\epsilon},\delta)$-DP such that $\tilde{\epsilon}\leq\epsilon$.
Property (i) essentially guarantees differential privacy (Dwork et al. 2006; Dwork and Roth 2014) at each training iteration. Property (ii) extracts the iteration-wise optimization, which expects that the difference measurement $\mathcal{F}_{\mathsf{Diff}}$ between the noisy model and pure model are small as possible. In other words, at each training iteration, the algorithm ensure $(\tilde{\epsilon},\delta)$-DP, while simultaneously seeking to minimize the divergence from an ideal output under a constrained privacy budget Given a fixed privacy budget $\epsilon/T$, improving utility expects to reduce the difference between $\mathcal{M}^{(t)}$ and non-noisy $Z^{\ast}$. Property (iii) makes sure that no additional privacy leakages are incurred in $\mathsf{DP\text{-}Hero}$ under privacy composition, which is the same as the standard DP guarantee. Overall, Definition 7 formalizes a utility-optimized perspective on differential privacy by requiring mechanisms to preserve per-iteration privacy guarantees while explicitly minimizing a utility loss objective relative to an optimal target.
Overview of DP Heterogeneous SGD
Before constructing DP heterogeneous SGD ($\mathsf{DP\text{-}Hero}$ SGD), we adopt the notations of DP-SGD by revisiting standard DP-SGD (Abadi et al. 2016). DP-SGD trains a model with parameters $\mathbf{W}$ by minimizing the empirical loss function $\mathcal{L}(\mathbf{W})$. For a random example $x_{i}$, DP-SGD computes gradients $\mathbf{g}(x_{i})\leftarrow\nabla(\mathbf{W},x_{i})$ with clipping value $C$, and adds noise to $\sum_{i}\mathbf{g}(x_{i})$ sampled from Gaussian distribution $\mathcal{N}(0,\sigma^{2}C^{2}\mathbf{I})$. An adversary cannot view the training process except for the DP-trained model.